127 grams of copper completely reacts with silver nitrate in the reaction Cu + 2 AgNO3 -> CuNO3)2 + 2 Ag

how many moles of AgNO3 are consumed?


1
Expert's answer
2021-06-09T12:54:03-0400

Molar Mass of copper = 63.546

127/63.546 = 1.9985 M

Molar Mass of AgNO3 = 169.87

Mole ratio = 1Cu : 2 AgNO3

= 2 × 1.9985 = 3.997 moles of AgNO3